# SiCP: Simultaneous Individual and Cooperative Perception for 3D Object Detection in Connected and Automated Vehicles

### v1

@@ 4 Experiments
Inference. At the inference stage, we use the average precision (AP) metric to evaluate the performance of all models on two OPV2V test sets (Default and Culver). The evaluation involved employing IoU thresholds of 0.5 and 0.7, respectively.
@@
Baselines. To facilitate comparison, we construct two baseline models dedicated to cooperative and individual perception. In the cooperative perception task, late fusion aggregates prediction outcomes from multiple agents, employing Non-Maximum Suppression (NMS) to produce the final results. Furthermore, we evaluate individual perception results against state-of-the-art standalone 3D detection networks such as PointPillars [13], PIXOR [28], SECOND [27], and VoxelNet [31]. Notably, SiCP diverges from established benchmarks where vehicles must initially project their individual point clouds onto the ego vehicle’s viewpoint; instead, our implementation mandates each vehicle to exclusively process data from its own perspective, mirroring real-world data-sharing scenarios among connected vehicles.
4.1 Quantitative Evaluations
@@
In the comparative analysis, we assess our model against existing benchmarks from two distinct perspectives: individual perception and cooperative perception. Our findings reveal that SiCP outperforms all other cooperative solutions, meanwhile, SiCP demonstrates satisfactory performance for individual perception scenarios.
@@
Cooperative Perception. Regarding cooperative perception, as shown in Table 1, SiCP achieves an impressive 85.72% AP with an IoU of 0.5, surpassing the second-best solution (AttFuse) which offers an 82.75% AP. Compared to the benchmark F-Cooper, SiCP exhibits a notable 8.2% enhancement. This improvement is attributed to the incorporation of the complementary feature fusion, which carefully merges two feature maps by learning optimized weights for every position within the overlapping feature map. The weights assigned to the ego vehicle and sender vehicle are mutually complementary, aligning seamlessly with the inherent logic of fusing feature maps.
Table 1: Comparison of cooperative perception performance
Default Culver
Methods AP@0.5 AP@0.7 AP@0.5 AP@0.7
Late Fusion 75.09 69.79 66.92 60.99
F-Cooper[2] 77.52 62.95 57.62 44.54
AttFuse[26] 82.75 72.63 72.02 59.76
V2X-ViT[25] 82.14 65.81 75.16 55.81
Where2comm[10] 78.72 64.84 75.84 59.04
SiCP (Ours) 85.72 73.64 77.49 62.42
Individual Perception. Traditional cooperative solutions like F-Cooper [2], AttFuse [25], Where2comm [10], and V2X-ViT [25], are designed to tackle cooperative perception challenges and do not consider individual perception at all. Therefore, when they are applied to process the data generated solely by the ego vehicle, their performance declines significantly. As shown in Table 2, the performance of V2X-ViT drops by 17.03%, compared to the standalone model PointPillars when the IoU is set at 0.7. This underscores the significance of considering individual perception in the design of cooperative perception models. Our developed SiCP solution showcases a significant improvement, with its performance closely approaching that of the best standalone models. The strength of our model stems from its specialized dedicated pipeline for individual perception, utilizing features derived solely from the ego vehicle.
Table 2: Comparison of individual perception performance
Default Culver
Methods AP@0.5 AP@0.7 AP@0.5 AP@0.7
PointPillars [13] 73.14 60.40 69.57 57.11
PIXOR [28] 61.34 40.04 48.84 33.48
SECOND [27] 72.85 62.63 70.53 59.32
VolxNet [31] 66.41 52.23 66.80 50.72
F-Cooper [2] 67.89 55.88 63.73 52.36
AttFuse [26] 65.67 53.44 64.13 51.60
V2X-ViT [25] 57.27 43.37 64.81 47.16
Where2comm [10] 57.79 43.76 66.51 51.49
SiCP (Ours) 73.14 61.22 69.67 56.53
4.2 DP-Net is a Lightweight Plug-and-Play Module
@@
We conduct a comparative analysis involving two pairs of comparisons, each comprising an original backbone model and the one extended by integrating DP-Net. The selected standalone 3D detection backbones are two representative network architectures in the field of 3D object detection: PointPillars [13] and VoxelNet [31]. As demonstrated in Table 3, DP-Net consistently demonstrates improvement across all baseline 3D detection backbones. These findings underscore the generic applicability of DP-Net, suggesting its potential to be integrated into other 3D object detection frameworks. Moreover, the proposed DP-Net is a lightweight solution as it does not significantly increase the number of parameters requiring training, i.e., DP-Net consists of only $0.13M$ parameters. If the underlying backbone is PointPillars, with a set of $7.27M$ parameters to be trained, the introduction of DP-Net increases the overall parameters by a mere $1.7\%$. With its lightweight design, SiCP achieves a latency of just $37.84ms$.
@@
Table 3: Existing 3D detection backbones can be extended by integrating DP-Net to address cooperative perception.
Default Culver
Backbones Methods AP@0.5 AP@0.7 AP@0.5 AP@0.7
PointPillars [13] IND 73.14 60.40 69.57 57.11
+DP-Net 85.72 73.64 77.49 62.42
VoxelNet [31] IND 66.41 52.23 66.80 50.72
+DP-Net 83.44 72.67 78.71 70.91
Figure 6: Illustrations of false positive (yellow) and false negative (blue) detection results for the individual perception scenarios.
Figure 7: Illustrations of false positive (yellow) and false negative (blue) detection results for the cooperative perception scenarios.
4.3 Qualitative Evaluations
@@
More qualitative results for individual perception are presented in Figure 6. Specifically, Figure 6 (a) showcases detection errors for F-Cooper, (b) illustrates errors for AttFuse, (c) displays errors for V2X-ViT, and (d) demonstrates errors of SiCP. The figure highlights instances of false positive and false negative detection results, attributed to incorrect feature extraction in existing cooperative solutions methods while attempting to address the individual perception task. In contrast, the SiCP method incorporates both individual and cooperative perception during feature extraction, leading to more precise object detection results.
@@
Regarding cooperative perception, as depicted in Figure 7, our proposed SiCP model also exhibits significantly fewer false negative and false positive detection results in comparison to (a) F-Cooper, (b) AttFuse, and (c) V2X-ViT. Notably, we observe a higher incidence of false negative detections in cooperative perception, compared to the individual perception scenarios. This is because any erroneous fusion can generate features that fail to indicate objects accurately, thereby causing false negatives. The superior performance of SiCP is attributed to DP-Net, which effectively fuses features and generates a more precise representation of objects within the feature maps.
4.4 Ablation Study
Impact of Complementary Fusion. To better understand the DP-Net, particularly the importance of complementary feature fusion, we exclude the branch responsible for applying the complementary weights (CW) to the received feature maps. As demonstrated in Table 4, the incorporation of the CW map results in a noteworthy enhancement in performance metrics, both at AP@0.5 and AP@0.7. This underscores the importance of weighting feature maps in a complementary fashion, emphasizing its pivotal role in augmenting the accuracy of object detection.
Table 4: Importance of complementary weights (CW)
Default Culver
Method AP@0.5 AP@0.7 AP@0.5 AP@0.7
w/o $CW$ 84.76 70.16 78.32 61.70
w. $CW$ 85.72 73.64 77.49 62.42
Impact of 1x1 Convolutional Layer. In the complementary feature fusion architecture, the stacked BEV feature maps need to be dimensionally reduced from 512 channels to 1 channel. We opt for a 1x1 convolutional layer for this purpose, facilitating feature interaction across channels. Although alternative methods exist, our experiments in Table 5 reveal that 1x1 convolutions mostly outperform other alternatives.
Table 5: Impact of the 1x1 convolutional layer.
Default Culver
Method AP@0.5 AP@0.7 AP@0.5 AP@0.7
Mean 85.31 70.64 77.45 61.91
Max 85.16 71.54 77.87 61.89
Ours 85.72 73.64 77.49 62.42
Impact of Number of Convolutional Layers. We incorporate two convolutional layers within the complementary feature fusion to learn the weight map. As shown in Table 6, increasing the convolutional layer number from 1 to 2 results in a substantial improvement at both AP@0.5 and AP@0.7. This finding indicates that deeper feature extraction capabilities contribute to enhancing object detection performance. However, further increasing the number to 3 led to a marginal decline in performance, suggesting the presence of a saturation point specific to our task. Therefore, we select the default configuration of 2 layers with 3x3 convolutions to strike a balance between complexity and performance.
Table 6: Impact of different numbers of convolutional layers.
# of Default Culver
layers AP@0.5 AP@0.7 AP@0.5 AP@0.7
1 85.37 72.38 76.74 61.54
2 85.72 73.64 77.49 62.42
3 84.84 70.58 77.83 61.01
Impact of Kernel Size. Our results, as presented in Table 7, indicate that the 3x3 convolutional kernel consistently delivers the best overall performance. This finding indicates the popularity of the 3x3 kernel size in modern neural networks, attributed to its balanced combination of computational efficiency and feature-capturing capabilities.
Table 7: Impact of kernel sizes.
Kernel Default Culver
size AP@0.5 AP@0.7 AP@0.5 AP@0.7
1x1 84.67 70.74 77.45 61.73
3x3 85.72 73.64 77.49 62.42
5x5 85.82 73.35 76.71 62.05
5 Conclusions

### v2

@@ IV Experiments
Inference. At the inference stage, following [29, 5], we use the Average Precision (AP) metric to evaluate the performance of all models on V2V4Real testset and OPV2V testsets (Default and Culver). The evaluation used Intersection over Union (IoU) thresholds of 0.5 and 0.7, respectively.
@@
Baselines. We construct baseline models specifically for individual and cooperative perception for comparison with the proposed SiCP method. In the individual perception task, the baseline model is no fusion, employing the advanced standalone 3D detection network PointPillars [1], alongside several SOTA deep fusion models: F-Cooper [4], AttFuse [5], V2X-ViT [6] and CoBEVT [7]. In the cooperative perception task, the baseline model is late fusion, combining final prediction outputs from multiple vehicles at a later stage, and the same deep fusion models as in the individual perception task. Notably, SiCP differs from established deep fusion baseline models, where vehicles must first project their respective point clouds onto the ego vehicle’s viewpoint; instead, our implementation requires each vehicle to process the data exclusively from its own perspective, reflecting real-world data sharing scenarios between connected vehicles.
IV-A Quantitative Evaluations
@@
In the comparative analysis, we evaluate our model against existing benchmarks on two datasets from two different perspectives: individual perception and cooperative perception. Our findings reveal that SiCP outperforms other cooperative solutions, meanwhile, SiCP demonstrates satisfactory performance for individual perception task.
Evaluation on V2V4Real dataset. In cooperative perception scenarios, our SiCP method demonstrates exceptional performance on the V2V4Real Dataset, as depicted in Table I. In a synchronization setting, where vehicles share perception data instantaneously, SiCP dominates with an impressive AP at an IoU=0.7. Notably, it surpasses CoBEVT [7] by a significant margin of 9.5% in the critical 0-30m range at this IoU, showcasing its strong, accurate, and reliable near-range detection capabilities crucial for ensuring safety in dense traffic environments. This improvement is attributed to the incorporation of the complementary feature fusion,
TABLE I: Individual and Cooperative Perception Evaluation on V2V4Real Dataset.
Method Individual Perception (AP@IoU=0.5/0.7)
overall 0-30m 30-50m 50-100m
F-CooperNo $\,{}_{\text{No Fusion}}$ [4] 35.4/17.6 58.0/29.1 23.1/12.3 5.6/2.8
AttFuseNo $\,{}_{\text{No Fusion}}$ [5] 30.2/15.1 51.2/25.9 19.8/10.7 4.4/2.2
V2X-ViTNo $\,{}_{\text{No Fusion}}$ [6] 35.3/16.5 57.0/26.3 22.2/11.5 5.1/2.7
CoBEVTNo $\,{}_{\text{No Fusion}}$ [7] 37.7/20.5 58.1/34.2 22.4/10.8 4.9/2.7
PointPillars [1] 38.6/23.3 62.5/38.9 25.3/14.9 5.7/3.2
SiCPNo $\,{}_{\text{No Fusion}}$(Ours) 38.0/22.3 60.4/37.9 24.2/12.4 4.1/2.0
Method Cooperative Perception (AP@IoU=0.5/0.7)
overall 0-30m 30-50m 50-100m
Late Fusion 42.2/24.1 67.4/32.6 25.3/10.9 11.7/6.7
F-Cooper [4] 47.7/20.2 67.9/31.7 31.0/13.6 23.5/7.2
AttFuse [5] 43.0/18.9 61.5/30.3 29.1/12.5 16.9/5.3
V2X-ViT [6] 49.2/20.2 70.7/30.3 31.0/14.2 26.9/8.5
CoBEVT [7] 51.0/23.0 68.4/34.2 34.9/15.1 26.7/8.8
SiCP (Ours) 50.0/25.6 71.6/43.7 35.8/16.0 26.5/8.9
@@
which carefully merges two feature maps by learning optimized weights for every position within the overlapping feature map. The weights assigned to the ego vehicle and sender vehicle are mutually complementary, aligning seamlessly with the inherent logic of fusing feature maps.
In individual perception scenarios, where vehicles operate without shared data from other vehicles, SiCP still maintains a strong presence, particularly in near and mid-range detection. Our SiCP solution showcases a significant improvement, with its performance closely approaching that of the advanced standalone model PointPillars [1].
Evaluation on OPV2V dataset. As shown in Table II, our SiCP surpasses competing methods with a standout 71.89% AP at IoU=0.7, reflecting a 1.75% improvement over the second-ranked V2X-ViT [6] in cooperative perception scenarios. Particularly in the Culver test set, SiCP secures a 63.02% AP at IoU=0.7, outpacing V2X-ViT by a substantial 3.26% thanks to its innovative approach in fusing and interpreting shared data among vehicles.
In individual perception scenarios, SiCP also exhibits strong performance, closely matching the top standalone model, PointPillars [1]. However, in the Default test set, CoBEVT [7] and AttFuse [5] drops by 4.97% and 7%, Moreover, in the Culver test set, CoBEVT [7] sees a further drop of 8.1% at IoU=0.7. This underscores the significance of considering individual perception in the design of cooperative perception models. The strength of our model stems from its specialized dedicated pipeline for individual perception, utilizing features derived solely from the ego vehicle.
TABLE II: Individual and Cooperative Perception Evaluation on OPV2V Dataset.
Method Individual Perception (AP@IoU=0.5/0.7)
Default test set Culver test set
F-CooperNo $\,{}_{\text{No Fusion}}$ [4] 68.79 56.76 75.31 59.82
AttFuseNo $\,{}_{\text{No Fusion}}$ [5] 67.94 53.89 73.61 57.64
V2X-ViTNo $\,{}_{\text{No Fusion}}$ [6] 72.26 57.09 75.19 56.52
CoBEVTNo $\,{}_{\text{No Fusion}}$ [7] 70.36 55.92 70.99 54.11
PointPillars [1] 73.84 60.89 77.71 62.21
SiCPNo $\,{}_{\text{No Fusion}}$(Ours) 73.72 60.89 76.36 61.28
Method Cooperative Perception (AP@IoU=0.5/0.7)
Default test set Culver test set
Late Fusion 77.03 62.95 75.59 58.45
F-Cooper [4] 79.71 63.95 68.26 48.86
AttFuse [5] 81.91 67.00 78.00 58.11
V2X-ViT [6] 85.62 70.14 79.09 59.76
CoBEVT [7] 84.77 68.11 75.22 55.42
SiCP (Ours) 85.64 71.89 79.10 63.02
IV-B DP-Net is a Robust Module on Alignment Error
The DP-Net effectively handles alignment errors arising from asynchronous communication and inaccurate localization between vehicles. As shown in Figure 5 (a), adapting to an asynchronous environment with a 100 ms communication delay, the resilience of SiCP becomes evident. In the 0-30m range, it maintains robust performance at IoU=0.7 with an AP of 41%, solidifying its ability to process delayed data efficiently. SiCP’s AP exceeds that of its closest competitor by 8.2%, underscoring its exceptional skill in near-range detection despite the challenges posed by asynchronous communication. Furthermore, SiCP demonstrates exceptional robustness to real-world challenges, specifically localization errors between vehicles. Through experiments that introduce Gaussian noise to simulate the errors (the noise is set to x, y location and yaw angle), as illustrated in Figure 5 (b), SiCP consistently maintains its superior performance. This showcases its ability to effectively manage inaccuracies in vehicle positioning, underscoring its practical utility and robustness in cooperative perception scenarios.
(a) Asynchronous on V2V4Real
(b) Loc. error on OPV2V
Figure 5: Robust response to asynchronous mode and localization error. SiCP outperforms other SOTAs across both datasets with IoU=0.7.
IV-C DP-Net is a Lightweight Plug-and-Play Module
@@
We conduct a comparative analysis involving two pairs of comparisons, each comprising an original backbone model and the one extended by integrating DP-Net. The selected standalone 3D detection backbones are two representative network architectures in the field of 3D object detection: PointPillars [1] and VoxelNet [3]. As demonstrated in Table III, DP-Net consistently demonstrates improvement across all baseline 3D detection backbones. These findings underscore the generic applicability of DP-Net, suggesting its potential to be integrated into other 3D object detection frameworks. Moreover, the proposed DP-Net is a lightweight solution as it does not significantly increase the number of parameters requiring training, i.e., DP-Net consists of only $0.13M$ parameters. If the underlying backbone is PointPillars, with a set of $7.27M$ parameters to be trained,
@@
TABLE III: Existing 3D detection backbones can be extended with DP-Net to address cooperative perception.
Backbone Method Default AP@IoU Culver AP@IoU
0.5 0.7 0.5 0.7
PointPillars [1] - 73.84 60.89 77.71 62.21
+DP-Net 85.64 71.89 79.10 63.02
VoxelNet [3] - 72.99 62.13 71.86 59.95
+DP-Net 83.23 69.36 77.38 62.73
the introduction of DP-Net increases the overall parameters by a mere $1.7\%$. With its lightweight design, SiCP achieves a latency of just $37.84ms$.
IV-D Qualitative Evaluations
@@
More qualitative results for individual perception and cooperative perception are presented in both datasets, shown in Figure 6 and Figure 7. Specifically, (a) showcases detection errors for F-Cooper, (b) illustrates errors for V2X-ViT, (c) displays errors for CoBEVT, and (d) demonstrates errors of SiCP. Regarding individual perception, both figures highlights instances of false positive and false negative detection results, attributed to incorrect feature extraction in existing cooperative solutions methods while attempting to address the individual perception task. In contrast, the SiCP method incorporates both individual and cooperative perception during feature extraction, leading to more precise object detection results.
@@
Regarding cooperative perception, our SiCP model also exhibits significantly fewer false negative and false positive detection results in comparison to (a) F-Cooper, (b) V2X-ViT, and (c) CoBEVT. Notably, we observe a higher incidence of false negative detections in cooperative perception, compared to the individual perception. This is because any erroneous fusion can generate features that fail to indicate objects accurately, thereby causing false negatives. The superior performance of SiCP is attributed to DP-Net, which effectively fuses features and generates a more precise representation of objects within the feature maps.
Figure 6: Illustrations of false positive and false negative in individual and cooperative perception on the OPV2V dataset.
Figure 7: Illustrations of false positive and false negative in individual and cooperative perception on the V2V4Real dataset.
IV-E Ablation Study
To better understand how SiCP can simultaneously handle individual and cooperative perception tasks, we conducted ablation studies on the individual perception pipeline and the complementary weight map. As evident from the results in Table IV, the individual perception pipeline is crucial for ensuring SiCP’s capability in individual perception tasks. Meanwhile, the complementary weight map highlights the importance of weighting feature maps in a complementary manner, crucial for effective feature fusion. Together, these components contribute to SiCP’s robustness in addressing various perception challenges.
TABLE IV: Ablation Study on the Individual Perception Pipeline and Complementary Weight Map ($1-M$) within the Cooperative Perception Pipeline. Results are reported in AP@IoU=0.7 on OPV2V dataset.
Individual Complementary Individual Cooperative
Perception Pipeline Weight Map Default Default
31.14 72.00
61.01 69.41
63.02 71.89
V Conclusions
@@
VI Acknowledgments
The work is supported by the National Science Foundation grants CNS-2231519, OAC-2017564, and ECCS-2010332.